OpenStudy (zuha):

What are the solutions? 16x^2+9=0

OpenStudy (irishboy123):

they are imaginary \(x^2= - \dfrac{9}{16}\) on an argand diagram you will find 2 solutions 

OpenStudy (zuha):

@IrishBoy123 Is the argand diagram simply a graph or is it something else? I've never seen it or heard of it before..

OpenStudy (irishboy123):

your solution is \[x= \sqrt{ - \dfrac{9}{16}}\] i don't know what you're learning but taking the sqrt of a -ve number is a bit of a challenge. you'll get 2 answers. is this what you expected ?? complex numbers have 2 squares roots, 3 cube roots, 4 quartic roots, all symmetrically arranged around the complex plane origin. if this is jibberish to you, you should say. perhaps there is a typo in what we are talking about :(

OpenStudy (zuha):

@IrishBoy123 Hm idk, this is a screen shot of the questions.

OpenStudy (zuha):

And these are the options, and like it says, the first one is correct (I took a quiz and I did bad on it so I'm reviewing the questions I messed up on on OpenStudy so I can be prepared to answer questions like them in the future lol)

OpenStudy (sshayer):

\[x^2=-\frac{ 9 }{ 16 }=\frac{ 9 }{ 16 }\times -1=\frac{ 9 }{ 16 }\iota^2\] \[x^2-\frac{ 9 }{ 16 }\iota^2=0\] \[x^2-\left( \frac{ 3 }{ 4 }\iota \right)^2=0\] \[\left( x+\frac{ 3 }{ 4 }\iota \right)\left( x-\frac{ 3 }{ 4 } \iota \right)=0\] x=?

OpenStudy (zuha):

@sshayer 3/4i, -3/4i?

OpenStudy (sshayer):

correct.

OpenStudy (zuha):

@sshayer Thank you for your help :)